OpenStudy (anonymous):

What is log 100000 = 5 in exponential form?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\large x=log_b~ y~~~~\rightarrow~~~y=b^x\]

OpenStudy (anonymous):

\[ \log 100000 = 5 \rightarrow 5 × 10^0\]

OpenStudy (anonymous):

(any \(log\) whose base is not specified has a natural base of \(10\)) therefore, \[\large 5=log_{10}~100000~~~~\rightarrow~~~~100000=10^5\]

OpenStudy (anonymous):

and that is \(true \) :)